All of these are necessary for perfect competition, EXCEPT: O differentiated products. O no barriers to market entry or exit.One of the innovations that helped globalization was: o the development of currency controls. O an increase in market demand.(Table) Based on the table, what is the average total product when Johns Salsa Company hires three people? Labor Total ProduFarmer Jean sells corn in a perfectly competitive market. The market price for a bushel of corn is $6. Jean has six hundred bIf there is no output level where price can cover average variable costs, the firm: O O will earn a normal profit by remaininWhich characteristic is an assumption in a perfectly competitive market structure? O O O O significant barriers to entry and
Answer #1

1.A competitive firm does not produce differentiated products.

Answer-differentiated products.

2.An increase in global rise in per capita income which further raised aggregate demand played a major role in globalisation.

Answer-Second option.

3.Average product=Total product/Output

AP=33/3

AP=11

Answer-11

4.Total cost=Fixed cost+Variable cost

TC=3,300+300

TC=3,600

Average cost=TC/Q

AC=3,600/600

AC=6

Because Price is equal to average cost,the producer is earning normal profit.

Answer-Last option.

5.A firm shuts down if it can't cover average variable cost in the short run.

Answer-Third option.

6.Perfect information exists between sellers and buyers in a competitive market structure.

Answer-Last option